# Compositional Coordination for Multi‑Robot Teams with Large Language Models

### v1

@@ Abstract
Multi-robot coordination has traditionally relied on a task-specific and expert-driven pipeline, where natural language mission descriptions are manually translated by domain experts into mathematical formulation, algorithm design, and executable code. This conventional process is labor-intensive, inaccessible to non-experts, and inflexible to changes in mission requirements. Here, we propose LAN2CB (Language to Collective Behavior), a novel framework that leverages large language models (LLMs) to streamline and generalize the multi-robot coordination pipeline. LAN2CB directly converts natural language mission descriptions into executable Python code for multi-robot systems through two key components: (1) Mission Decomposition for Task Representation, which parses the mission into a task graph with dependencies, and (2) Code Generation, which uses the task graph and a structured knowledge base to generate deployable robot control code. We further introduce a dataset of natural language mission specifications to support development and benchmarking. Experimental results in both simulation and real-world settings show that LAN2CB enables effective and flexible multi-robot coordination from natural language, significantly reducing the need for manual engineering while supporting generalization across mission types. Website: https://sites.google.com/view/lan2cb
@@ I Introduction
Problem: Multi-robot coordination for complex collective behaviors, including coverage, formation, foraging, and exploration, has been extensively studied for decades [1, 2, 3, 4, 5, 6, 7]. However, most prior coordination methods target a specific and narrowly defined coordination task. This highlights the need for a general and unified framework that can address a wide range of coordination problems in a systematic way. The conventional multi-robot development pipeline starts with a natural language (NL) problem description, enabling both expert and non-expert users to specify objectives and constraints. An expert translates this description into a mathematical problem (typically an optimization problem). This inevitably involves making some assumptions. Following this, a domain expert designs algorithms to solve the formulated problem, which are subsequently implemented and validated on robotic hardware. However, this pipeline has three limitations: It is labor-intensive, inaccessible to non-experts, and inflexible. The first two limitations are apparent due to the significant use of skilled experts, despite the repetitive nature of problem formulation, algorithm design, and code implementation. The third becomes apparent when small changes to the problem statement or new requirements emerge, and the entire pipeline has to be manually redesigned.
@@
Background: LLMs are trained on massive amounts of text data to understand and generate human-like language, allowing them to interpret natural language prompts, perform contextual reasoning, and produce coherent responses or actions in both visual and linguistic tasks[13, 14, 15, 16, 17, 18, 19]. In robotics, LLMs are increasingly being used to improve high-level decision-making, task planning, and human-robot interaction [20, 21, 22, 23, 24, 25, 26, 27, 28]. Applications include interpreting natural language instructions, generating structured plans, assisting in code generation for robot control, and enabling dialogue-based coordination among multi-robot systems. The key advantages of LLMs in robotics are their flexibility, adaptability to new tasks without extensive retraining, and ability to bridge the gap between human intent and machine execution, significantly reducing the need for manual programming and domain-specific engineering.
@@
Contributions: 1. We propose, implement, and thoroughly evaluate an LLM-based framework, Language to Collective Behavior (LAN2CB), which addresses the limitations of the traditional pipeline and directly turns natural language into executable Python code for robot teams. LAN2CB consists of two main modules (Fig. 1). The first module is Mission Decomposition for Task Representation, which takes the mission specified in natural language as input, identifies the necessary tasks to be completed, and analyzes the dependency between tasks. The output of the first module, all ready-to-run nodes from a task graph, is given to the second module Code Generation, together with a preconstructed knowledge base, to generate executable code for robots. 2. We design a dataset of natural language mission specifications for multi-robot coordination. 3. Our experiments, both in simulation and hardware, demonstrate that LAN2CB produces effective multi-robot coordination from natural language. They also show that LAN2CB is flexible and extendable, obviating the need to design from scratch whenever the mission changes and providing a standard format to incorporate existing knowledge into its knowledge base.
@@ II-B1 Single-Robot Integration
Recent work has explored integrating large language models (LLMs) with single robots to enable more intuitive human-robot interaction and high-level task planning. These approaches leverage the language understanding and reasoning capabilities of LLMs to interpret natural language commands and translate them into executable actions. Because natural language is inherently tied to contexts and semantics, LLMs are particularly effective in understanding tasks such as navigation, exploration, and instruction following. In these scenarios, LLMs can help robots ground language inputs into spatial or semantic representations that guide behavior in complex, partially known environments [23, 24, 22]. In addition, real-time residency [20] and anomaly detection [25] have been integrated into LLM-based robotic pipelines to improve robustness in dynamic and unforeseen situations.
@@ II-B2 Team-Level Integration
Extending LLMs to multi-robot systems introduces challenges in coordination, communication, and scalability, with research still in the early stages exploring their use for flexible, robust multi-robot coordination [39]. Some works use LLMs to translate natural language into formal representations [40, 41, 42, 43], but they are limited to simple missions and cannot handle multi-robot teams or decomposable long-horizon tasks. Others propose prompt-based frameworks for multi-robot coordination using LLMs [27, 28]. For instance, Zhao et al. introduced RoCO [27], where robots coordinate via interactive dialogues. However, such frameworks face limitations with large robot teams (e.g., $\geq 10$) and complex language-based missions. Shyam et al. proposed SMART-LLM [28], an LLM-based task planner for multi-robot teams with task decomposition and allocation, but it lacks domain knowledge, limits generalization beyond in-context examples, and cannot handle motion-level language commands. Our work is closely related to GenSwarm [44], which translates natural language missions into executable Python code. While both aim to enable language-driven multi-robot coordination, our framework differs in two key ways. First, our modular design supports a broader range of coordination problems, including multi-team tasks with interdependencies, unlike GenSwarm’s focus on simple, single-team scenarios. Second, we incorporate a structured knowledge base to guide the LLM, enhancing extensibility, whereas GenSwarm relies solely on in-context learning, limiting generality.
Figure 2: A comprehensive workflow demonstration of LAN2CB.   user- or expert-provided content.   LLM-generated mission descriptions.   LLM-generated code.
@@ III Problem Statement
    Constraints: Additional requirements can be imposed, such as entering a region from a specific direction, maintaining speed limits, or adhering to prescribed durations for actions.
@@
    Trigger: Tasks can include trigger conditions, enabling actions to be terminated or modified in response to environmental events (e.g., if a robot comes within a certain distance, say 1 m, of a forbidden region, it switches to an alternative goal).
@@
    Atomic Task Finish Condition: Each atomic task has a clearly defined completion criterion, such as reaching a designated position within a user-specified tolerance or herding objects within a given range of the target region.
@@ IV Method
    Mission Analysis: Given natural language mission descriptions from users, LLMs autonomously decompose missions into atomic task specifications and analyze intertask dependencies. For each task, LLMs extract robot assignments, relevant constraints, triggers, finish conditions, contextual hints, and other contextual requirements. Tasks that are immediately executable are then identified and forwarded to the next stage.
@@
    Code Generation: Based on the mission description, the system first generates the mission-level completion criteria, which are produced once for the entire mission cycle. For each selected task, considering its constraints, trigger conditions, finish conditions, and hints, the system separately generates: (a) robot execution code, (b) code for evaluating trigger conditions, and (c) code for assessing task completion.
@@
    Execution: The generated code is deployed to the robots for execution. During run-time, if any trigger condition or task completion criterion is met, the workflow transitions to stage 4. If the overall mission completion condition is satisfied, it proceeds directly to stage 5.
@@
    Mission Completion: Once all mission requirements are met, the workflow is terminated and the mission is formally marked as complete.
@@ IV-A Mission Analysis
Mission analysis can be divided into two phases: mission understanding and ready-to-run action node selection. However, user-provided natural language mission descriptions often lack explicit definitions of all atomic tasks and their execution dependencies. Consequently, directly prompting LLMs for general mission understanding may result in incomplete or ambiguous task specifications and overlook critical execution dependencies. To address these challenges, we first standardized mission descriptions by using an expert-provided template, and then decomposed the mission understanding stage into three steps: 1) dependency analysis, 2) comprehensive mission analysis, and 3) ready-to-run action node(s) selection.
Dependency Analysis: To initiate dependency analysis, we prompt an LLM to extract all atomic tasks from the mission description, initially disregarding trigger conditions and any tasks that may be activated post-trigger. The LLM then analyzes the dependency relationships among these atomic tasks. If triggers are activated during execution, we re-invoke the LLM to update the dependency analysis accordingly.
Comprehensive Mission Analysis: Building on the results of dependency analysis, we combine these atomic tasks and their dependencies with the standardized mission description to prompt the LLM for comprehensive mission analysis. The mission is represented in a behavior tree structure, with the overall mission as the root and atomic tasks as action nodes. Non-leaf nodes of type Parallel or Sequence capture task execution topologies according to dependencies. For each node, we instruct the LLM to generate a unique index, identify the node type, task name, status, constraints, trigger conditions, finish conditions, hints, and children. For action nodes, the LLM also specifies the action type as well as the associated robots, objects, regions, and points. See Tab. I for detailed field descriptions of each BehaviorTreeNode.
@@
Ready-to-run Action Node(s) Selection: After constructing the complete behavior tree and integrating the dependency analysis results, we ask LLMs to identify action nodes that are immediately executable. Specifically, these are nodes for which all dependencies are satisfied and all preconditions are met in the current environment state. The ready-to-run action nodes are then selected and forwarded for downstream processing, code generation. This approach ensures that robot actions are initiated only when all prerequisite tasks and constraints have been resolved, thereby supporting safe and efficient mission execution. After selecting action nodes that are ready-to-run, we instruct LLMs to update the status of all nodes in the behavior tree, marking the selected nodes and their parent nodes as Running.
@@ IV-B Code Generation
After mission analysis and the identification of ready-to-run action nodes, we prompt LLMs to generate Python code to control all robots in the environment. The code generation process consists of four key components, executed synchronously with the robots’ control loop (Fig. 1 and  2):
  1. 1.
    Robot execution code generation;
  2. 2.
    Trigger condition code generation;
  3. 3.
    Action node finish condition code generation;
  4. 4.
    Mission finish condition code generation.
Robot Execution Code Generation: Recent works have shown that LLMs can effectively assist code generation [45, 46, 47, 48]. However, to increase the success rate and avoid hallucination, we still need to provide enough prior information to LLMs instead of asking LLMs to generate all code from scratch [49]. To this end, we construct a meta behavior library for robot execution. To ensure flexibility and the capability to process multiple action nodes simultaneously, the meta-behavior library comprises three sub-libraries:
@@
Action Node Finish Condition Code Generation: For each selected action node, LLMs generate code to evaluate completion based on its specified finish condition. For example, the task of an action node is to visit three points. The action node finishes when all three points are visited.
Mission Finish Condition Code Generation: Mission finish condition generation is similar in nature to composite task finish condition generation. However, in this case, we only need to prompt LLMs to generate the mission finish condition code once for the entire mission.
@@ IV-C Execution
Once all code scripts are generated, the system begins or resumes execution. During execution, if the trigger condition is satisfied, we prompt LLMs to update the dependency analysis by removing the current task that caused the trigger and adding new tasks associated with this trigger. The updated dependency analysis, combined with the mission description, is then used to prompt LLMs to generate a new behavior tree. Afterward, LLMs select new ready-to-execute action nodes, and generate code for these nodes. If the finish condition of any action nodes is satisfied, LLMs update the behavior tree, including their parent nodes if the finish condition of these parent nodes solely relies on the completion of their children. LLMs then select new ready-to-run action nodes and generate corresponding code. If the finish condition of any composite nodes is satisfied, LLMs update the behavior tree by marking all its child nodes that have not yet started or are still running as Failure and marking the composite nodes as Success. The execution continues iteratively until the mission finish condition is fully satisfied.
Figure 3: Quantitative results for nine multi-robot missions. Missions #1–3 require LLMs to analyze simple missions and utilize the existing meta-behavior library for code generation. For example, mission #1 asks three robots on the left to visit the top left region, and three robots on the right to visit the three right regions separately. After these two tasks are completed, all robots then enclose the top left region. Missions #4-6 go further, requiring LLMs to synthesize code beyond the provided library (e.g., generating spiral trajectories). For example, mission #4 asks all robots to form an ’H’ first, followed by forming an ’I’. Since the library only supports point visits, LLMs need to compute each robot’s position. Missions #7-9 are the most complex, involving intricate dependencies, triggers, or adaptive behaviors. For example, mission #7 starts with asking the robots on the left to visit points next to the top green region and the three robots on the right to enclose the green region on the right. If any robot on the left comes within 1.5 units of the bottom blue region, a trigger redirects all three to the top-left region. Please check the website for more details.
@@ V Experiments
The experimental study is designed to rigorously assess LAN2CB under a variety of multi-step, long-horizon missions. We focus on three questions.
  1. 1)
    How well does LAN2CB perform across diverse missions?
  2. 2)
    How does the use of a standardized template for mission descriptions affect the performance of LAN2CB compared to using raw natural language?
  3. 3)
    Can LAN2CB be reliably deployed on real robots?
@@ V-A Performance Across Diverse Missions
To the best of our knowledge, no standard dataset exists for evaluating multi-step, long-horizon multi-robot mission execution under natural language instructions. Therefore, we curate a comprehensive mission suite comprising three groups (nine missions total), each designed to test the distinct capabilities of our system.
@@
  • Category C: Advanced Missions. These missions demand advanced analysis and complex code generation. They may involve intricate dependencies, trigger conditions, or non-trivial mission-level finish criteria, such as region coverage requirements, dynamic triggers, or adaptive behaviors. (See Missions #7, #8, and #9 in Fig. 3.)
Evaluation Metrics: We quantitatively evaluate missions using the metrics: (1) Avg. Success Rate: The proportion of missions completed successfully as intended, without human intervention. (2) Avg. Token Number and Cost: The total number of LLM tokens used per mission and related cost. (3) Avg. Tokens per Error: The average number of tokens consumed per error given the output tokens of LLMs.
@@
Results: We evaluate all missions using GPT-4.1, conducting each mission five times. Detailed results are provided in Tab. II. By standardizing raw mission descriptions from users with an expert-provided template, LAN2CB achieves a 100% success rate on Category A (basic missions), a 93% success rate on Category B (basic missions with advanced code generation) and 87% success rate on Category C (advanced missions). In addition, the system demonstrates strong robustness, achieving approximately 0.12 million tokens per error across all missions.
These results demonstrate that LAN2CB reliably analyzes mission requirements, including task sequencing, extraction of essential information from natural language instructions, selection of ready-to-run nodes, and dynamic updates to behavior tree-style representations. Notably, the high success rate in Category B and Category C highlights its ability to perform advanced code generation tasks. These include synthesizing robot trajectories not natively supported by the behavior library, generating custom formations, creating complex execution conditions such as specifying approach directions or target-following durations, generating triggers, and generating mission finish conditions. Overall, our results indicate that LAN2CB is both robust and adaptable for diverse multi-robot mission execution scenarios.
Figure 4: Snapshots of real-world experiments.   task not activated.   task running.   task finished.

### v2

@@ Abstract
Multi-robot coordination has traditionally relied on a mission-specific and expert-driven pipeline, where natural language mission descriptions are manually translated by domain experts into mathematical formulation, algorithm design, and executable code. This conventional process is labor-intensive, inaccessible to non-experts, and inflexible to changes in mission requirements. Here, we propose LAN2CB (Language to Collective Behavior), a novel framework that leverages large language models (LLMs) to streamline and generalize the multi-robot coordination pipeline. LAN2CB transforms natural language (NL) mission descriptions into executable Python code for multi-robot systems through two core modules: (1) Mission Analysis, which parses mission descriptions into behavior trees, and (2) Code Generation, which leverages the behavior tree and a structured knowledge base to generate robot control code. We further introduce a dataset of natural language mission descriptions to support development and benchmarking. Experiments in both simulation and real-world environments demonstrate that LAN2CB enables robust and flexible multi-robot coordination from natural language, significantly reducing manual engineering effort and supporting broad generalization across diverse mission types. Website: https://sites.google.com/view/lan-cb
@@ I Introduction
Problem: Multi-robot coordination for complex collective behaviors, including coverage, formation, foraging, and exploration, has been extensively studied for decades [1, 2, 3, 4, 5, 6, 7]. However, most prior coordination methods target a specific and narrowly defined coordination task. This highlights the need for a general and unified framework that can address a wide range of coordination problems in a systematic way. The conventional multi-robot development pipeline starts with a natural language (NL) problem description, enabling both expert and non-expert users to specify objectives and constraints. An expert translates this description into a mathematical formulation, typically as an optimization problem, which inevitably requires additional assumptions. Following this, a domain expert designs algorithms to solve the formulated problem, which are subsequently implemented and validated on robotic hardware. However, this pipeline has three limitations: It is labor-intensive, inaccessible to non-experts, and inflexible. The first two limitations are apparent due to the significant use of skilled experts, despite the repetitive nature of problem formulation, algorithm design, and code implementation. The third becomes apparent when even minor changes to the problem statement or new requirements emerge, as the entire pipeline must be manually redesigned.
@@
Background: LLMs are trained on massive amounts of text data to understand and generate human-like language, allowing them to interpret natural language prompts, perform contextual reasoning, and produce coherent responses or actions in both visual and linguistic tasks[13, 14, 15, 16, 17, 18, 19]. In robotics, LLMs are increasingly being used to improve high-level decision-making, task planning, and human-robot interaction [20, 21, 22, 23, 24, 25, 26, 27, 28]. Applications include interpreting natural language instructions, generating structured plans, assisting in code generation for robot control, and enabling dialogue-based coordination among multi-robot systems. The key advantages of LLMs in robotics include their flexibility, ability to adapt to new tasks without extensive retraining, and capacity to bridge the gap between human intent and machine execution, thereby significantly reducing the need for manual programming and domain-specific engineering.
@@
Contributions: 1. We propose, implement, and thoroughly evaluate an LLM-based framework, Language to Collective Behavior (LAN2CB), which addresses the limitations of the traditional pipeline and turns natural language mission descriptions into executable plans and Python code for robot teams. LAN2CB consists of two main modules (Fig. 1). The first module is Mission Analysis, which takes the mission specified in natural language as input, identifies the necessary tasks to be completed, and analyzes the dependency between tasks. The output of the first module, all ready-to-run nodes from a behavior tree, is given to the second module Code Generation, together with a pre-constructed knowledge base, to generate executable code for robots. 2. We design a dataset of natural language mission specifications for multi-robot coordination. 3. Our experiments, both in simulation and hardware, demonstrate that LAN2CB produces effective multi-robot coordination from natural language. They also show that LAN2CB is flexible and extendable, obviating the need to design from scratch whenever the mission changes and providing a standard format to incorporate existing knowledge into its knowledge base.
@@ II-B1 Single-Robot Integration
Recent work has explored integrating LLMs with single robots to enable more intuitive human-robot interaction and high-level task planning. These approaches leverage the language understanding and reasoning capabilities of LLMs to interpret natural language commands and translate them into executable actions. Because natural language is inherently tied to contexts and semantics, LLMs are particularly effective in understanding tasks such as navigation, exploration, and instruction following. In these scenarios, LLMs can help robots ground language inputs into spatial or semantic representations that guide behavior in complex, partially known environments [23, 24, 22]. In addition, real-time resilience [20] and anomaly detection [25] have been integrated into LLM-based robotic pipelines to improve robustness in dynamic and unforeseen situations.
@@ II-B2 Team-Level Integration
Extending LLMs to multi-robot systems introduces challenges in coordination, communication, and scalability, with research still in the early stages exploring their use for flexible, robust multi-robot coordination [39]. Some works use LLMs to translate natural language into formal representations [40, 41, 42, 43], but they are limited to simple missions and cannot handle multi-robot teams or decomposable long-horizon tasks. Others propose prompt-based frameworks for multi-robot coordination using LLMs [27, 28]. For instance, Zhao et al. introduced RoCO [27], where robots coordinate via interactive dialogues. However, such frameworks face limitations with large robot teams (e.g., $\geq 10$) and complex language-based missions. Shyam et al. proposed SMART-LLM [28], an LLM-based task planner for multi-robot teams with task decomposition and allocation, but it lacks domain knowledge, limits generalization beyond in-context examples, and cannot handle motion-level language commands. Our work is closely related to GenSwarm [44], which translates natural language task descriptions into executable Python code. While both aim to enable language-driven multi-robot coordination, our framework differs in two key ways. First, our modular design supports a broader range of coordination problems, including multi-team tasks with interdependencies, unlike GenSwarm’s focus on simple, single-team scenarios. Second, we incorporate a structured knowledge base to guide the LLM, enhancing extensibility, whereas GenSwarm relies solely on in-context learning, limiting generality.
Figure 2: A comprehensive demonstration of LAN2CB.   user- or expert-provided content.   LLM-generated mission descriptions.   LLM-generated code.
@@ III Problem Statement
    Constraints: Additional requirements can be imposed, such as entering a region from a specific direction or adhering to prescribed durations for actions.
@@
    Trigger: Tasks can include trigger conditions that allow actions to be terminated or modified in response to environmental events. For example, if a robot comes within a specified distance (e.g., 1 meter) of a forbidden region, it needs to switch to another task.
@@
    Task Finish Condition: Each atomic task has a clearly defined completion criterion, such as reaching a designated position within a user-specified tolerance or herding objects within a given range of the target region.
@@ IV Method
    Mission Analysis: Given natural language mission descriptions from users, the LLM autonomously decomposes missions into atomic task specifications and analyze inter-task dependencies. For each task, the LLM extracts robot assignments, relevant constraints, triggers, finish conditions, contextual hints, and other contextual requirements. Tasks that are immediately executable are then identified and forwarded to the next stage.
@@
    Code Generation: Based on the mission description, the system first generates the mission finish condition, which is generated once for the entire mission cycle. For each selected task, considering its constraints, trigger conditions, finish conditions, and hints, the system separately generates code for: (a) robot execution, (b) trigger conditions, and (c) task finish conditions.
@@
    Execution: The generated code is deployed to the robots for execution. During run-time, if any trigger condition or task finish condition is met, the workflow transitions to stage 4. If the overall mission finish condition is satisfied, it proceeds directly to stage 5.
@@
    Mission Completion: Once the mission finish condition is met, the system terminates and the mission is formally marked as complete.
@@ IV-A Mission Analysis
Mission analysis consists of two key phases: mission understanding and ready-to-run action node selection. However, user-provided natural language mission descriptions often omit explicit definitions of atomic tasks and their execution dependencies. As a result, directly prompting LLMs for general mission understanding can yield incomplete or ambiguous task specifications and may overlook critical execution dependencies. To address these challenges, the mission understanding phase begins by standardizing the mission description using an expert-provided template. We then perform dependency analysis to explicitly extract task relationships, which serve as the basis for a comprehensive mission analysis. Afterward, the LLM selects ready-to-run action nodes based on the results of the dependency and comprehensive mission analyses.
Dependency Analysis: To initiate dependency analysis, we prompt the LLM to extract all atomic tasks from the mission description, initially disregarding trigger conditions and any tasks that may be activated post-trigger. The LLM then analyzes the dependency relationships among these atomic tasks. If triggers are activated during execution, we re-invoke the LLM to update the dependency analysis accordingly.
Comprehensive Mission Analysis: Building on the results of dependency analysis, we combine these atomic tasks and their dependencies with the standardized mission description to prompt the LLM for comprehensive mission analysis. The mission is represented in a behavior tree structure, with the overall mission as the root and atomic tasks as action nodes. Non-leaf nodes of type Parallel or Sequence capture task execution topologies according to dependencies. For each node, we instruct the LLM to generate a unique index, identify the node type, task name, status, constraints, trigger conditions, finish conditions, hints, and children. For action nodes, the LLM also specifies the action type as well as the associated robots, objects, regions, and points. See Tab. I for detailed field descriptions of the BehaviorTreeNode.
@@
Ready-to-Run Action Node(s) Selection: After constructing the complete behavior tree and integrating the dependency analysis results, we ask the LLM to identify action nodes that are immediately executable. Specifically, these are nodes for which all dependencies are satisfied and all preconditions are met in the current environment state. The ready-to-run action nodes are then selected and forwarded for downstream processing, code generation. This approach ensures that robot actions are initiated only when all prerequisite tasks and constraints have been resolved, thereby supporting safe and efficient mission execution. After selecting action nodes that are ready-to-run, we instruct the LLM to update the status of all nodes in the behavior tree, marking the selected nodes and their parent nodes as Running.
@@ IV-B Code Generation
After mission analysis and the identification of ready-to-run action nodes, we prompt the LLM to generate Python code to control all robots in the environment. The code generation process consists of four key components, executed synchronously with the robots’ control loop (Fig. 1 and  2): 1) Robot execution code generation; 2) Trigger condition code generation; 3) Action node finish condition code generation; 4) Mission finish condition code generation.
Robot Execution Code Generation: Recent works have shown that LLMs can effectively assist code generation [45, 46, 47, 48]. However, to improve success rates and mitigate hallucinations, it is more effective to provide core functions in advance and expose them as APIs, rather than asking the LLM to generate all code from scratch [49]. To this end, we construct a meta behavior library for robot execution. To ensure flexibility and the capability to process multiple action nodes simultaneously, the meta-behavior library comprises three sub-libraries:
@@
Action Node Finish Condition Code Generation: For each selected action node, the LLM generates code to evaluate whether the node’s specified finish condition has been met. For example, if the task is to visit three points, the action node is considered complete when all three points have been visited.
Mission Finish Condition Code Generation: Mission finish condition generation is similar in nature to composite task finish condition generation. However, the LLM only needs to generate the mission finish condition code once for the entire mission.
@@ IV-C Execution
Once all code scripts are generated, the system initiates or resumes execution. During execution, if a trigger condition is satisfied, we prompt the LLM to update the dependency analysis by removing the current task that caused the trigger and adding any new tasks associated with the trigger. The updated dependency analysis, together with the mission description, is then used to prompt the LLM to generate a new behavior tree. The LLM subsequently selects new ready-to-run action nodes and generates code for them. If the finish condition of any action node is satisfied, the LLM updates the behavior tree. This may also include updating parent nodes if their completion depends solely on the completion of their children. The LLM then selects new ready-to-run action nodes and generates the corresponding code. If the finish condition of any composite node is satisfied, the LLM updates the behavior tree by marking all of its child nodes that have not started or are still running as Failure, and marking the composite node as Success. Execution proceeds iteratively until the mission finish condition is satisfied.
Figure 3: Quantitative results for nine multi-robot missions. Please check the website for more details.
@@ V Experiments
The experimental study is designed to rigorously evaluate LAN2CB under a variety of multi-step, long-horizon missions. We focus on three questions: 1) How effectively does LAN2CB perform across diverse missions? 2) How does the use of a standardized template for mission descriptions impact the performance of LAN2CB compared to raw natural language input? 3) Can LAN2CB be reliably deployed on real-world robotic platforms?
@@ V-A Performance Across Diverse Missions
To the best of our knowledge, no standard dataset exists for evaluating multi-step, long-horizon multi-robot mission execution under natural language mission descriptions. Therefore, we curate a comprehensive mission suite comprising three groups (nine missions total), each designed to test the distinct capabilities of our system.
@@
  • Category C: Advanced Missions. These missions demand advanced analysis and complex code generation. They may involve intricate dependencies, trigger conditions, or non-trivial mission finish conditions, such as region coverage requirements, dynamic triggers, or adaptive behaviors. (See Missions #7, #8, and #9 in Fig. 3.)
Evaluation Metrics: We quantitatively evaluate missions using the metrics: (1) Avg. Success Rate: The proportion of missions completed successfully as intended, without human intervention. (2) Avg. Token Number and Cost: The total number of LLM tokens used per mission and related cost. (3) Avg. Tokens per Error: The average number of tokens consumed per error given the output tokens of the LLM.
@@
Results: We evaluate all missions using GPT-4.1, conducting each mission five times. Detailed results are provided in Tab. II. By standardizing raw mission descriptions from users with an expert-provided template, LAN2CB achieves a 100% success rate on Category A (basic missions), a 93% success rate on Category B (basic missions with advanced code generation) and 87% success rate on Category C (advanced missions). In addition, the system demonstrates strong robustness, achieving $\sim 0.12$ million tokens per error across all missions. These results demonstrate that LAN2CB reliably analyzes mission descriptions, including identifying task dependencies, extracting essential information, selecting ready-to-run nodes, and dynamically updating behavior trees. Notably, the high success rates in Category B and Category C highlight its ability to perform advanced code generation tasks. These tasks include synthesizing robot trajectories not natively supported by the behavior library, generating custom formations, creating complex execution conditions such as specifying approach directions or target-following duration, and generating both triggers and mission finish conditions. Overall, our results indicate that LAN2CB is both robust and adaptable for diverse multi-robot mission execution scenarios.
Figure 4: Snapshots of real-world experiments.   idle,   running,   finished.